OpenStudy (kakarot):

What is the value of d? 2d – 4 = 4 a.)d=1 b.)d=7 c.)d=4 d.)d=3

OpenStudy (texaschic101):

2d - 4 = 4 --- add 4 to both sides 2d - 4 + 4 = 4 + 4 -- simplify 2d = 8 -- divide both sides by 2 d = 8/2 d = 4 questions ?

OpenStudy (kakarot):

no

OpenStudy (anonymous):

Mhm nice.

OpenStudy (anonymous):

2d-4=4 2d=0

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

d=4

OpenStudy (pawanyadav):

OK check by putting d=4

OpenStudy (kakarot):

d=4 is correct. thanks